# Toward Universal Laws of Outlier Propagation

### v2

@@ Abstract
We argue that Algorithmic Information Theory (AIT) admits a principled way to quantify outliers in terms of so-called randomness deficiency. For the probability distribution generated by a causal Bayesian network, we show that the randomness deficiency of the joint state decomposes into randomness deficiencies of each causal mechanism, subject to the Independence of Mechanisms Principle. Accordingly, anomalous joint observations can be quantitatively attributed to their root causes, i.e., the mechanisms that behaved anomalously. As an extension of Levin’s law of randomness conservation, we show that weak outliers cannot cause strong ones when Independence of Mechanisms holds. We show how these information theoretic laws provide a better understanding of the behaviour of outliers defined with respect to existing scores.
@@ 1 Introduction
Anomaly detection plays a crucial role in business, technology, and medicine. Typical use cases range from fraud detection in finance and online trading [Donoho, 2004], performance drops in manufacturing lines [Susto et al., 2017] or cloud computing applications [Gan et al., 2021, Ma et al., 2020, Hardt et al., 2023], health monitoring in intensive care units [Maslove et al., 2016], to explaining extreme weather and climate events [Zscheischler et al., 2022]. It has motivated a vast effort towards developing methodologies relevant to outlier analysis. As example, we refer the reader to some of the early works in statistics and computer science [Freeman, 1995, Rocke and Woodruff, 1996, Rousseeuw and Leroy, 2003, Aggarwal, 2017]. In complex systems, an anomaly will typically cause a large cascades of related anomalies [Panjei et al., 2022]. In order to mitigate them, it is not sufficient to merely detect the anomalies; we must also identify which of the anomalies was the root cause [Budhathoki et al., 2022, Ikram et al., 2022, Li et al., 2022, Hardt et al., 2023, Wang et al., 2023b, a]. Thus, we implicitly face the counterfactual question of what conditions could have been different to prevent the (usually undesired) anomalous event.
@@
Causal Bayesian networks offer a framework that supports both kinds of modular description, specifying causal relations via a directed acyclic graph (DAG) $G$ with random variables $X_{1},\dots,X_{n}$ as nodes [Pearl, 2009, Spirtes et al., 1993]. Due to the causal Markov condition [Pearl, 2009], the joint distribution factorizes according to
$P(X_{1},\dots,X_{n})=\prod_{i=1}^{n}P(X_{j}\mid\operatorname{{PA}}_{j}),$ (1)
where $\operatorname{{PA}}_{j}$ denotes the parents of $X_{j}$ in $G$, i.e., its direct causes. We will think of each conditional distribution $P(X_{j}\mid\operatorname{{PA}}_{j})$ as an independent mechanism of the system, which can in principle be changed or replaced without changing the others (see 2.1 and 2.2 in Peters et al. [2017] for a historic overview). Following this modular view, our root cause analysis will assume that in the case of an anomalous observation, most of the mechanisms worked as expected; thus, the anomaly can be blamed on a small number of mechanisms that act as root causes, in alignment with Schölkopf et al. [2021]’s “sparse mechanism shift hypothesis”.
To our knowledge, Budhathoki et al. [2022] provide the most elaborate formalization of the idea of attributing anomalies to mechanisms. We develop our concepts starting from this baseline.
@@ 1.1 Outlier scores from p-values
To quantitatively attribute an anomalous event to upstream nodes, Budhathoki et al. [2022] first introduce what they call an Information Theoretic (IT) outlier score via
@@
where $x$ denotes an observation of the random variable $X$, and $\tau:{\cal X}\to\mathbb{R}$ is an appropriate feature statistic, whose choice will be discussed later [^1] While Budhathoki et al. [2022] use the natural base $e$, we use base $2$ logarithms to align with binary program lengths in algorithmic information theory. In effect, we express the outlier score in units of bits [Frank, 2005]..
$\lambda_{\tau}$ can be viewed as a statistical test of the null hypothesis that $x$ was sampled from $P$: setting the base of the logarithm to $2$ yields the p-value $2^{-\lambda_{\tau}(x)}$. A small p-value (or large $\lambda_{\tau}$) corresponds to an unusual sample under $P$, which can be labeled an outlier. Since $x$ is a single observation, anomaly scoring thus reduces to classical hypothesis testing on a sample size of $1$ [Shen, 2020, Vovk, 2020] [^2] Note that Tartakovsky et al. [2012] discuss anomalies as change points over multiple observations. However, we focus on anomalies confined to an individual observation..
1.2 Quantitative root cause analysis in causal Bayesian networks
To quantitatively attribute an anomalous observation $(x_{1},\dots,x_{n})$ to different mechanisms, where each $x_{j}$ represents a value of the corresponding variable $X_{j}$, the arXiv version of Budhathoki et al. [2022] define conditional outlier scores:
@@ 1.2 Quantitative root cause analysis in causal Bayesian networks
They demonstrate that this score can be equivalently interpreted as measuring the outlierness of the noise term [^3] This aligns in spirit with Von Kügelgen et al. [2023], where events are also backtracked by attributing them to the noise variables.. The feature functions $\tau_{j}$ can be node-specific, which is essential when the variables $X_{j}$ have different characteristics (e.g., different dimensionality or data types). When the specific choice of $\tau$ is not crucial for the discussion, we will simplify notation by dropping the subscript. The arXiv version of Budhathoki et al. [2022] extends this framework by introducing a joint outlier score through “convolution”:
@@
-- $\log\sum_{i=1}^{n-1}\frac{(\sum_{j=1}^{n}\lambda(x_{j}\mid\operatorname{{pa}}_{j}))^{i}}{i!}.\vspace{-0.2cm}$
The second term serves as a correction factor that ensures the score maintains the properties of an IT score, provided that the conditional distributions have densities with respect to the Lebesgue measure.
@@ 1.3 Monotonicity of scores
Okati et al. [2024] formalize an intuitive principle about anomaly propagation: unless the connecting mechanisms themselves are also anomalous, a moderate outlier in a cause (measured by $\lambda(x_{1})$) should not produce an extreme outlier in its effect (measured by $\lambda(x_{2})$). This can be demonstrated for the bivariate causal relation $X_{1}\to X_{2}$ as follows:
@@
In other words, whenever we generate an anomaly $x_{1}$ randomly from $P(X_{1}\mid\lambda(X_{1})\geq c)$ and then generate $X_{2}$ via its mechanism $P(X_{2}\mid X_{1})$, the resulting outlier score is unlikely to be much larger than $c$. The intuitive reason is that the probability occupied by events with high scores is much smaller than that of a given event having a lower score.
Note that there is no statement for any single $x_{1}$ – some of them may generate outliers that are higher with large probability – but “most” of them do not. The reason that we cannot make any statement about particular values is crucial for the present paper, because it shows how independence of mechanisms comes into play: whenever one chooses, on purpose, an $x_{1}$ such that it results in a large outlier downstream, there is no reason that this particular $x_{1}$ needs to have high outlier score. One can easily think, for instance, of the deterministic relation $X_{2}=f(X_{1})$, where $f$ is a highly non-linear function that results in rather anomalous values $x_{2}$ for some tiny region of values $x_{1}$ which are not anomalous.
1.4 Limitations of current approach
@@
$\lambda_{\tau}$ defines outliers in terms of a feature statistic $\tau$ that is chosen in advance. Sometimes, we only know what makes a sample abnormal after seeing it, so we would like to be able to choose $\tau$ with hindsight. For example, suppose $P$ is univariate Gaussian. It seems natural to define $\tau(x):=|x-\mu|$, so that the anomaly detector flags extremely high or low values. However, we might also like to flag observations such as $x=0$ or $x=\mu$: while such $x$ may lie in a region of high density, the probability of obtaining such a specific value is zero under $P$. Although any other value has also probability zero, one may agree that observing $x=0$ and $x=\mu$ would be surprising because these values are special (in a sense specified later).
As another example, suppose each component of a multivariate observation $x$ is the reading of a different sensor. The signal transmission from all $d$ sensors may be broken in such a way that all components equal to the same constant “idle” state $c$. Such coincidences also indicate an unusual event.
Aggarwal [2016] describe a broad variety of different outliers beyond the above toy examples: These can be unusual frequencies of words in text documents [Mohotti and Nayak, 2020], unexpected patterns in images [Yakovlev et al., 2021], unusually large cliques in graphs [Hooi et al., 2016], or points lying in low density regions [Breunig et al., 2000].
@@
While (1.2) nicely decomposes the joint outlier score into mechanism-specific conditional scores, a key limitation is that this decomposition relies on defining the joint score based on a sum of conditional scores. This does not imply that any reasonable outlier score (e.g., using a generic feature function $\tau$) for the joint observation can be decomposed in this manner.
@@ 1.5 Our contributions
The contribution of this paper is purely conceptual. We do not propose another outlier detection or root cause analysis method, but instead provide a theoretical framework for calibrating and interpreting outlier scores. The framework ensures that outlier scores meet three critical criteria: (i) Comparability across diverse probability spaces and data modalities. (ii) Non-increasingness along causal chains of downstream effects, regardless of variable modalities. (iii) Well-defined attribution of joint system outlier scores to anomalies of mechanisms.
@@
Good information theoretic concepts enable many nice theorems, but they are often hard to work with in practice. However, together with distributional assumptions (e.g. Gaussianity) they can boil down to simple concepts (e.g. linear algebraic expressions). The resulting formulae may be valid and useful beyond the distributional assumptions (e.g. by virtue of linear algebra).
@@
The paper is structured as follows. Section 2 describes concepts from statistics, information theory, and causality that we build upon. Section 3 derives the decomposition, and Section 4 shows that the joint score is non-increasing under marginalization. Section 5 discusses simple examples and Section 6 a toy experiment. For a cleaner exposition, we defer some formal proofs and definitions to the supplementary material.
@@ 2.1 Statistical testing with e-values instead of p-values
While p-values are the most famous measure of evidence in statistical testing, e-values are recently gaining popularity for their superior ability to aggregate evidence across multiple tests [Ramdas and Wang, 2024]. These values are inconsistently scaled in the literature, with e-values being comparable to reciprocals of p-values and exponentials of algorithmic randomness scores. To remove any obfuscation coming from scaling conventions, we introduce the following definitions.
@@
Similarly, a statistic $\lambda:\mathcal{X}\rightarrow[-\infty,\infty]$ is a p-test (or e-test) in log form, if $2^{\lambda}$ is a p-test (or e-test) in ratio form. Our convention is to use lowercase symbols like $\lambda$ to indicate log form.
By Markov’s inequality, every e-test is also a p-test. Conversely, Vovk and Wang [2021] describe a number of ways to calibrate any given p-test into an e-test. For more details on tests, see Section 8 in the supplementary material.
p-tests provide a straightforward way to control the Type I error rate, i.e., false positives. To ensure that samples from $P$ are labeled as anomalies at a rate no larger than a desired threshold $\epsilon$, we flag only those samples whose p-test scores are above $1/\epsilon$ (i.e., below $\epsilon$ when expressed in probability form).
Since every e-test is also a p-test, e-tests also achieve this false positive rate; however, they are more conservative. In return, e-tests offer many advantages related to composability and optional stopping [Grünwald et al., 2020, Ramdas et al., 2023, Ramdas and Wang, 2024], and we will find them more convenient for decomposing anomaly scores by mechanism.
@@ 2.2 Basic notions from algorithmic information theory
The intuition behind universal tests is that an observation is anomalous precisely when it is more compressible than usual for the underlying random process. To formalize description lengths, we fix a universal prefix-free Turing machine. The conditional Kolmogorov complexity $K(x\mid y)$ is the bit length of the shortest program $p$ that outputs $x$ when given access to another string $y$ [Li and Vitányi, 2019]. It satisfies the Kraft inequality $\sum_{x}2^{-K(x\mid y)}<1$. When $y$ is an empty string, we write $K(x)$. Just as Shannon’s entropy measures information content for a probability distribution, $K(x)$ measures it for an individual sample $x$.
@@
Here, $z^{*}$ denote a shortest program that outputs $z$, and $\stackrel{{\scriptstyle+}}{{=}}$ denotes equality up to a constant dependent on the universal machine, but not on $x$ or $y$ [^4] Note that $x^{*}$ contains more information than $x$, because $x$ is easily generated from $x^{*}$ but not vice versa. Conditioning on $x$ instead of $x^{*}$ would result in error terms that are often constant, and at worst logarithmic in the length of $x$ [Gács, 2021].. We say $x$ and $y$ are conditionally independent, given $z$, if $I(x:y\mid z^{*})\stackrel{{\scriptstyle+}}{{=}}0$.
@@ 2.3 Universal tests
A Martin-Löf or Levin test (i.e., semicomputable p-test or e-test, respectively) can be thought of as combining all computable anomaly scoring algorithms, in order to detect the broadest possible variety of anomalies. Section 9 in the supplementary material contains full details; here we give a brief overview.
@@
The domination property provides a natural way to compare the power of different statistical tests. The class of semicomputable tests contains a universal test that dominates all the others:
@@
is a universal e-test, dominating all other semicomputable e-tests.
The intuition behind Eq. 4 is that typical samples from a distribution $P$ are optimally compressed by encodings of length $-\log P(x)$. When a sample $x$ can be compressed beyond this theoretical limit, we consider it “anomalous”. This includes the case where we observe $x=0$ from a discretized centered Gaussian distribution. Despite being the mode of the distribution, 00 is considered anomalous because its negative log-likelihood $-\log P(x)$ is substantially larger than its Kolmogorov complexity $K(0)\stackrel{{\scriptstyle+}}{{=}}0$.
Note that when a distributional estimate $\hat{P}$ is inferred from data, the “reconstruction loss” $-\log\hat{P}(x)$ can be used as an outlier score [Bishop, 1993]. Without further considerations, this score is not calibrated because there may be a huge region with low density, and thus all low density points together may still be likely. However, if the term $K(x\mid P^{*})$ is also small, we know that $x$ is special within this huge set. [^5] Note that conformal p-values [Bates et al., 2021] do not require any parametric assumptions or density estimation, since their coverage guarantees rely on exchangeability alone. However, statements on conditional coverage [Barber et al., 2019] are too weak for our purpose of calibrating conditional outlier scores.
@@
2.4 Algorithmic Markov condition and independence of mechanisms
Our goal is to demonstrate how the randomness deficiency decomposes across causal mechanisms in a Bayesian network, subject to postulates that connect algorithmic information theory to causality. To achieve this, we establish a theoretical framework based on fundamental postulates.
Janzing and Schölkopf [2010] propose an adaptation of (1) that characterizes algorithmic dependencies between individual objects, rather than statistical dependencies between random variables:

### v3

@@ Abstract
When a variety of anomalous features motivate flagging different samples as outliers, Algorithmic Information Theory (AIT) offers a principled way to unify them in terms of a sample’s randomness deficiency. Subject to the algorithmic Markov condition on a causal Bayesian network, we show that the randomness deficiency of a joint sample decomposes into a sum of randomness deficiencies at each causal mechanism. Consequently, anomalous observations can be attributed to their root causes, i.e., the mechanisms that behaved anomalously. As an extension of Levin’s law of randomness conservation, we show that weak outliers cannot cause strong ones. We show how these information theoretic laws clarify our understanding of outlier detection and attribution, in the context of more specialized outlier scores from prior literature.
@@ 1 Introduction
Generically speaking, an outlier is an anomalous or atypical event, suggesting possible interference and/or downstream effects. Anomaly detection plays a crucial role in business, technology, and medicine. Typical use cases range from fraud detection in finance and online trading [Donoho, 2004], performance drops in manufacturing lines [Susto et al., 2017] and cloud computing applications [Gan et al., 2021, Ma et al., 2020, Hardt et al., 2023], health monitoring in intensive care units [Maslove et al., 2016], to explaining extreme weather and climate events [Zscheischler et al., 2022]. It has motivated a vast effort towards developing methodologies relevant to outlier analysis. For examples, we refer the reader to some of the early works in statistics and computer science [Freeman, 1995, Rocke and Woodruff, 1996, Rousseeuw and Leroy, 2003, Aggarwal, 2017].
In complex systems, an anomaly may cause a large cascade of related anomalies [Panjei et al., 2022]. To mitigate them, it is not enough to merely detect anomalies; we must also identify which of the anomalies was the root cause [Budhathoki et al., 2022, Ikram et al., 2022, Li et al., 2022, Hardt et al., 2023, Wang et al., 2023b, a]. Thus, we implicitly face the counterfactual question of what conditions could have been different to prevent the (usually undesired) anomalous event.
@@
Causal Bayesian networks offer such a modular description, specifying causal relations via a directed acyclic graph (DAG) $G$ with random variables $X_{1},\dots,X_{n}$ as nodes [Pearl, 2009, Spirtes et al., 1993]. Under the causal Markov condition [Pearl, 2009], the joint distribution factorizes according to
$P(X_{1},\dots,X_{n})=\prod_{j=1}^{n}P(X_{j}\mid\operatorname{{PA}}_{j}),$ (1)
where $\operatorname{{PA}}_{j}$ denotes the parents of $X_{j}$ in $G$, i.e., its direct causes. We will think of each conditional distribution $P(X_{j}\mid\operatorname{{PA}}_{j})$ as an independent mechanism of the system, which can in principle be changed or replaced without changing the others (see 2.1 and 2.2 in Peters et al. [2017] for a historical overview).
Provided that most of the mechanisms operate normally, we can blame an anomalous joint observation $x_{1},\dots,x_{n}$ on a small number of mechanisms that we call root causes, in alignment with Schölkopf et al. [2021]’s “sparse mechanism shift hypothesis”. Budhathoki et al. [2022] formalize additional concepts for attributing anomalies to mechanisms, providing a starting point for our theoretical development.
@@ 1.1 Outlier scores from p-values
Budhathoki et al. [2022] introduce what they call an Information Theoretic (IT) outlier score [^1] While Budhathoki et al. [2022] use the natural base $e$, we use base $2$ logarithms to align with binary program lengths in algorithmic information theory. In effect, we express the outlier score in units of bits [Frank, 2005]. via
@@
where $x$ denotes an observation of the random variable $X$, and $\tau:{\cal X}\to\mathbb{R}$ is a feature statistic whose highest values we consider anomalous.
$\lambda_{\tau}$ can be viewed as a statistical test of the null hypothesis that $x$ was sampled from $P$: setting the base of the logarithm to $2$ yields the p-value $2^{-\lambda_{\tau}(x)}$. A small p-value (or large $\lambda_{\tau}$) corresponds to an unusual sample under $P$, which can be labeled an outlier. Since $x$ is a single observation, anomaly scoring thus reduces to classical hypothesis testing on a sample size of $1$ [Shen, 2020, Vovk, 2020] [^2] Note that Tartakovsky et al. [2012] discuss anomalies as change points over multiple observations, whereas we focus on anomalies confined to an individual observation..
1.2 Quantitative root cause analysis in causal networks
In order to quantify the contribution of different upstream mechanisms to an anomalous observation, the arXiv version of Budhathoki et al. [2022] defines the conditional outlier scores:
@@ 1.2 Quantitative root cause analysis in causal networks
They demonstrate that this score can be interpreted as measuring the anomalousness of the mechanism’s noise term, [^3] This aligns in spirit with Von Kügelgen et al. [2023] where events are also backtracked by attributing them to the noise variables. whenever $\lambda_{\tau}(X_{j}\mid PA_{j})$ is independent of $PA_{j}$. The feature functions $\tau_{j}$ can be node-specific, which is essential when the variables $X_{j}$ have different characteristics (e.g., different dimensionality or data types). When the specific choice of $\tau$ is not crucial for the discussion, we will simplify notation by dropping the subscript. The arXiv version of Budhathoki et al. [2022] extends this framework by introducing a joint outlier score through “convolution”:
@@
-- $\log\sum_{i=1}^{n-1}\frac{(\sum_{j=1}^{n}\lambda(x_{j}\mid\operatorname{{pa}}_{j}))^{i}}{i!}.$
The second term serves as a correction to ensure the score maintains the properties of an IT score, provided that the conditional distributions have densities with respect to the Lebesgue measure.
@@ 1.3 Monotonicity of scores
Okati et al. [2024] formalize an intuitive principle of anomaly propagation: unless the connecting mechanisms also behave anomalously, a moderate outlier (measured by $\lambda(x_{1})$) should not cause an extreme outlier (measured by $\lambda(x_{2})$). This can be stated for the bivariate causal relation $X_{1}\to X_{2}$ as follows:
@@
In other words, whenever we generate an anomaly $x_{1}$ randomly from $P(X_{1}\mid\lambda(X_{1})\geq c)$ and then generate $X_{2}$ via its mechanism $P(X_{2}\mid X_{1})$, the resulting outlier score is unlikely to be much larger than $c$. The reason is that the probability of an event of the form $\lambda(X_{j})\geq c$ decreases exponentially in $c$.
Note that Lemma 1.1 says nothing about any individual $x_{1}$ – some values of $x_{1}$ may generate stronger outliers with high probability, although “most” do not. For instance, it’s easy to construct a deterministic relation $X_{2}=f(X_{1})$, whose non-linearity maps a tiny region of non-anomalous values $x_{1}$ to anomalous values $x_{2}$. On the other hand, whenever this occurs, it must either be the case that the nonlinearity of $f$ is specially “tuned” to pick out those specific $x_{1}$; or else, $x_{1}$ is anomalous in some other sense not captured by the feature $\tau_{1}$. The algorithmic Markov condition will rule out the former case, prompting the need for a general outlier score to capture the latter case.
1.4 Limitations of current approaches
@@
$\lambda_{\tau}$ defines outliers in terms of a feature statistic $\tau$ that is chosen in advance. Sometimes, we only know what makes a sample abnormal after seeing it, so we would like to be able to choose $\tau$ with hindsight. For example, suppose $P$ is univariate Gaussian. It seems natural to define $\tau(x):=|x-\mu|$, so that the anomaly detector flags extremely high or low values. However, we might also like to flag observations such as $x=0$ or $x=\mu$, as it would be a surprising coincidence to see precisely these values. While the probability density at these points may be high, intuitively it seems much more likely to guess these specific numbers as a result of some alternative process different from $P$, than it is to sample them from $P$ (which indeed occurs with probability zero).
As another example, suppose each component of a continuous multivariate observation $x$ represents the reading of a different sensor. The signal transmission from all $d$ sensors may break in such a way that all components give the same constant “idle” state $c$. Such coincidences also indicate an unusual event.
Aggarwal [2016] describe a broad variety of different outliers beyond the above toy examples: These can be unusual frequencies of words in text documents [Mohotti and Nayak, 2020], unexpected patterns in images [Yakovlev et al., 2021], unusually large cliques in graphs [Hooi et al., 2016], or points lying in low-density regions [Breunig et al., 2000].
@@
While (1.2) nicely decomposes the joint outlier score into mechanism-specific conditional scores, a key limitation is that this decomposition relies on defining the joint score based on a sum of conditional scores. This does not suggest that any reasonable outlier score (e.g., using a generic feature function $\tau$) for the joint observation can be decomposed in this manner.
@@ 1.5 Our contributions
This paper’s contributions are purely conceptual. We do not propose a practical algorithm for outlier detection or root cause analysis, as our core definitions can only be computed in the infinite-runtime limit. Instead, we provide a theoretical framework for calibrating and interpreting outlier scores. The framework ensures that outlier scores meet three critical criteria: (i) Comparability across diverse probability spaces and data modalities. (ii) Non-increasingness along causal chains of downstream effects, regardless of variable modalities. (iii) Well-defined attribution of joint system outlier scores to anomalies of mechanisms.
@@
Good information-theoretic concepts enable many nice theorems, but they are often hard to work with in practice. However, together with distributional assumptions (e.g. Gaussianity) they can boil down to simple concepts (e.g. linear algebraic expressions). The resulting formulae may be valid and useful beyond the distributional assumptions (e.g. by virtue of linear algebra).
@@
The paper is structured as follows. Section 2 reviews concepts from statistics, information theory, and causality that we build upon. Section 3 derives the decomposition, and uses it to show that weak outliers do not cause stronger ones. Section 4 discusses simple examples, and Section 5 a toy experiment. For a cleaner exposition, we defer some formal proofs and definitions to the supplementary material.
@@ 2.1 Statistical testing with e-values instead of p-values
While p-values are the most famous measure of evidence in statistical testing, e-values are recently gaining popularity for their superior ability to aggregate evidence across multiple tests [Ramdas and Wang, 2024]. These values are inconsistently scaled in the literature, with e-values being comparable to reciprocals of p-values and exponentials of algorithmic randomness scores. To remove any obfuscation arising from scaling conventions, we introduce the following definitions.
@@
Similarly, a statistic $\lambda:\mathcal{X}\rightarrow[-\infty,\infty]$ is a p-test (or e-test) in log form, if $2^{\lambda}$ is a p-test (or e-test) in ratio form. Our convention is to use lowercase symbols like $\lambda$ to indicate log form. For more details on tests, see Section 7 in the supplementary material.
p-tests provide a straightforward way to control the Type I error rate, i.e., false positives. While we want to be flexible about what kinds of anomalies to consider, we should insist that most samples from $P$ are not outliers. To ensure that samples from $P$ are labeled as outliers at a rate no larger than a desired threshold $\epsilon$, we can choose a p-test and flag only those samples whose scores are above $1/\epsilon$ (i.e., below $\epsilon$ when expressed in probability form).
By Markov’s inequality, every e-test is also a p-test. Hence, e-tests achieve the same false positive rate; however, they are more conservative. In return, e-tests offer many advantages related to composability and optional stopping [Grünwald et al., 2020, Ramdas et al., 2023, Ramdas and Wang, 2024]. Our main result, Theorem 3.2, concerns a general-purpose e-test that conveniently decomposes into a sum of contributions, attributable to a causal network’s mechanisms.
@@ 2.2 Basic notions from algorithmic information theory
In a very general sense, an observation $x$ may be deemed an outlier whenever it belongs to a low-probability set that is flagged by some computable test. This condition can be rephrased in terms of the length of a compressed description of $x$.
To formalize description lengths, we fix a universal prefix-free Turing machine, and discretize our variables so that their values can be written as strings. The conditional Kolmogorov complexity $K(x\mid y)$ is the bit length of the shortest program $p$ that outputs $x$ when given access to another string $y$ [Li and Vitányi, 2019]. It satisfies the Kraft inequality $\sum_{x}2^{-K(x\mid y)}<1$. When $y$ is the empty string, we simply write $K(x)$. Just as Shannon’s entropy measures information content for a probability distribution, $K(x)$ measures it for an individual sample $x$.
@@
[…]
@@ 2.3 Universal tests
Semicomputable p- and e-tests are called Martin-Löf and Levin tests, respectively. Each class of tests can be aggregated, resulting in universal tests that effectively combine every conceivable anomaly scoring algorithm, in order to detect the broadest possible variety of outliers. Section 8 in the supplementary material contains more details; here we give a brief overview.
@@
The domination property provides a natural way to compare the power of different statistical tests. Since Definition 2.1 restricts us from increasing a test everywhere, it is perhaps surprising that the class of semicomputable tests contains a universal test that dominates all the others.
@@
is a universal Levin test, dominating all other Levin tests (i.e., semicomputable e-tests).
The intuition behind Eq. 4 is that typical samples from a distribution $P$ are optimally compressed by encodings of length $-\log P(x)$. When a sample $x$ can be compressed beyond this theoretical limit, we consider it anomalous. This includes the case where we observe $x=0$ from a discretized centered Gaussian distribution. Despite being the mode of the distribution, 00 is considered anomalous because its negative log-likelihood $-\log P(x)$ is substantially larger than its Kolmogorov complexity $K(0)\stackrel{{\scriptstyle+}}{{=}}0$.
@@
When a distributional estimate $\hat{P}$ is inferred from data, [Bishop, 1993] uses the “reconstruction loss” $-\log\hat{P}(x)$ as an outlier score. Without further considerations, this score is not calibrated, because there may be a huge region with low density, with high total probability. However, if the term $K(x\mid P^{*})$ is also small, then we can say that $x$ is special within this huge set. [^5] Note that conformal p-values [Bates et al., 2021] do not require any parametric assumptions or density estimation, since their coverage guarantees rely on exchangeability alone. However, statements on conditional coverage [Barber et al., 2019] are too weak for our purpose of calibrating conditional outlier scores.
2.4 Algorithmic Markov condition
In order to identify root causes of an anomaly, we would like to decompose the randomness deficiency into a sum of contributions from each causal mechanism in a Bayesian network. Our framework for doing so requires an algorithmic analogue of the causal Markov condition. Specifically, Janzing and Schölkopf [2010] propose an adaptation of (1) that characterizes algorithmic dependencies between individual objects, rather than statistical dependencies between random variables: